OpenStudy (anonymous):

Can anyone help me I need to solve this for d: S=n/2(2a+(n-1)d)

OpenStudy (anonymous):

I can't figure it out sorry. I tried.

OpenStudy (anonymous):

S = n/2 [ 2a+(n-1)d ] 2S/n = 2a+(n-1)d 2S/n - 2a = (n-1)d => d = ( 2S/n) - 2a ) / ( n -1)